40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 11 16:29:33 2013, Jenne, Update, Locking, Temp oplev for PR2; ITMX temporarily has no oplev 
    Reply  Tue Feb 12 00:50:08 2013, yuta, Update, Locking, PR2 oplev spectra PR2oplev.png
Message ID: 8066     Entry time: Tue Feb 12 00:50:08 2013     In reply to: 8058
Author: yuta 
Type: Update 
Category: Locking 
Subject: PR2 oplev spectra 

I wanted to see if PR2 motion makes PRC beam motion or not, using temporary oplev to PR2.
I could not measure the coherence between beam motion and PR2 motion, because I couldn't lock half-PRC today.
But I took spectra of PR2 oplev anyway.

Result:

  Below are the spectra of PR2 oplev outputs (taken using C1:SUS-ITMX_OL(PIT|YAW)_IN1). Bottom plot is POP DC during half-PRC locked yesterday.
PR2oplev.png

Discussion:
  We see bump in PR2 oplev output at ~ 2-3 Hz. But we cannot say this is a evidence for PR2 motion making PRC beam motion because no coherence measurement was done. Also, oplev might be just seeing the ITMX stack motion.

  Resonant frequency of TTs measured were at ~ 1.8-1.9 Hz (elog #8054), but we cannot clearly see these peaks in oplev outputs. Did resonant frequency shifted because of different damping condition?

ELOG V3.1.3-